Firm ABC has the following production function 𝑄 = 6𝐾 3 + 4𝐾𝐿 βˆ’ 𝐿 3 i. Determine the average products of capital (K) and labour (L). [4 Marks] ii. Find the marginal products of capital and labour (MPK and MPL)


1
Expert's answer
2022-01-24T08:39:24-0500

Solution:

i.). Average Product of Capital (APK) = "\frac{Q}{K}" = 6K3 + 4KL – L3 "\div" K = 6K2 + 4L – "\frac{L^{3} }{K}"


Average Product of Labor (APL) = "\frac{Q}{L}" = 6K3 + 4KL – L3 "\div" L = "\frac{6K^{3} }{L}" + 4K – L2

Β 

ii.). Marginal Product of Capital (MPK) = "\frac{\partial Q} {\partial K}" = 18K2 + 4L


Marginal Product of Labor (MPL) = "\frac{\partial Q} {\partial L}" = 4K – 3L2